OpenStudy (darkbluechocobo):

Find the missing value. y is directly proportional to x and inversely proportional to the square of z. y = 30 when x = 60 and z = 10. Find y when x = 2 and z = 2.

OpenStudy (perl):

y = k*x/z^2

OpenStudy (anonymous):

is this algebra II?

OpenStudy (darkbluechocobo):

It is precal. and 30=k*60/100?

OpenStudy (anonymous):

Dude... ur on ur own

OpenStudy (darkbluechocobo):

k=18

OpenStudy (darkbluechocobo):

@perl would this be y=9?

OpenStudy (darkbluechocobo):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

y directly proportional to x mean y=kx the other one @perl gave it to you

OpenStudy (darkbluechocobo):

wait so we are using direct and indirect?

OpenStudy (xapproachesinfinity):

no you have to use one equation i just divide the meaning of each perl gave you the equation above find k=? and answer the question for x=2 and z=2

OpenStudy (darkbluechocobo):

so 30=60/10^2